OpenStudy (anonymous):

Find the slope of a line passing through (6, 7) and (2, 9)

OpenStudy (tyteen4a03):

Remember the slope formula: \(S = \large{\frac{y_2 - y_1}{x_2 - x_1}}\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

So this is how you set it up: \[\frac{ 9-7 }{ 2-6 } = ?\]